# Gaussian curvature and Lyapunov exponent as probes of black hole phase transitions

### v1

@@ IV Critical Exponent from Lyapunov Exponents and Gaussian Curvature
Critical exponent characterize the behavior of thermodynamic systems near critical point. This section begins by briefly reviewing how the Lyapunov exponent $\lambda$ of black holes can be used to derive critical exponents for black hole phase transitions using a standard approach proposed in Refs. Banerjee and Roychowdhury (2012a, b). We subsequently examine whether the Gaussian curvature $K$ exhibits an analogous behavior.
Consider the expansion of $\lambda$ near the critical point $\tilde{r}_{c}$
$\lambda_{+}=\lambda_{c}+\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}(\tilde{r}_{+}-\tilde{r}_{c})+\mathcal{O}(\tilde{r}_{+}-\tilde{r}_{c})^{2},$ (38)
[…]
$\tilde{T}=\tilde{T}_{c}+\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}(\tilde{r}_{+}-\tilde{r}_{c})^{2}+\mathcal{O}(\tilde{r}_{+}-\tilde{r}_{c})^{3}.$ (39)
This leads to
$\tilde{r}_{+}-\tilde{r}_{c}\approx(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}.$ (40)
Substituting $\tilde{r}_{+}-\tilde{r}_{c}$ into the $\lambda$ expansion yields
$\lambda_{+}-\lambda_{c}=\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}.$ (41)
This implies the following relation:
$\Delta\lambda\sim(\tilde{T}-\tilde{T}_{c})^{\delta_{\lambda}},$ (42)
with the critical exponent $\delta_{\lambda}=1/2$. This result aligns with the analysis for non-spherically symmetric black holes in Refs. Chen et al. (2025); Yang et al. (2025a), and spherically symmetric black holes in Refs. Guo et al. (2022b); Shukla et al. (2024); Lyu et al. (2024). Following their work, the order parameter is defined as $\Delta\lambda\equiv\lambda_{l}-\lambda_{s}$, where $\lambda_{l}$ and $\lambda_{s}$ denote the Lyapunov exponents of large and small black holes, respectively.
We now investigate whether the Gaussian curvature $K$ exhibits similar behavior. We define $\Delta K$ as
$\Delta K\equiv|K(\tilde{r}_{+})|-|K(\tilde{r}_{c})|.$ (43)
Using the relation Eq. (21), we obtain
$\Delta K=\lambda_{+}^{2}-\lambda_{c}^{2}=(\lambda_{+}+\lambda_{c})(\lambda_{+}-\lambda_{c}).$ (44)
Substituting the expression Eq. (41) for $(\lambda_{+}-\lambda_{c})$ yields
$\Delta K=\left(\lambda_{+}+\lambda_{c}\right)\left(\frac{\partial\lambda}{\partial\tilde{r}_{+}}\right)_{c}\left[\frac{1}{2}\left(\frac{\partial^{2}\tilde{T}}{\partial\tilde{r}_{+}^{2}}\right)_{c}\right]^{-\frac{1}{2}}(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}.$ (45)
Since $(\lambda_{+}+\lambda_{c})$ is a sum of positive real numbers, i.e. a nonzero constant, we conclude
$\Delta K\sim(\tilde{T}-\tilde{T}_{c})^{\delta_{K}},$ (46)
with identical critical exponent $\delta_{K}=1/2$.
@@ V Discussion
In this work, we systematically investigate the profound connections between spacetime geometry, thermodynamics, and chaos dynamics during first-order phase transitions of black holes. We begin by reviewing the behavior of first-order phase transitions of black holes as thermodynamic systems, noting the limitations of conventional approaches that rely on thermodynamic potentials. We then introduce the Gaussian curvature $K$ as an intrinsic geometric quantity, and using a known relation $K(r_{LR})=-\lambda^{2}(r_{LR})$ for unstable null orbits, we link intrinsic geometry to dynamical chaos and black hole phase transitions.
@@
Through numerical analysis of three spherically symmetric black holes, the RN-AdS, the Hayward-AdS, and the Hayward-Letelier-AdS models, we find that when system parameters $\tilde{Q}$ and $\tilde{g}$ enter the spinodal region, the free energy exhibits a swallowtail structure, while both the Gaussian curvature $K$ and the Lyapunov exponent $\lambda$ display multivalued behavior. The temperature intervals over which this occurs align precisely with the thermodynamic spinodal region. Near the critical point, both $K$ and $\lambda$ diverge with a critical exponent of $1/2$, following the scaling relation $(\tilde{T}-\tilde{T}_{c})^{\frac{1}{2}}$, consistent with mean-field theory.
@@
Furthermore, we demonstrate that in the absence of a phase transition ($\tilde{g}>\tilde{g}_{c}$), the Gaussian curvature $K$ varies monotonically and the geometric multivaluedness disappears. These results provide the first evidence of spacetime degeneracy accompanying black hole phase transitions and indicate that Gaussian curvature can serve as a diagnostic tool for phase transitions independent of thermodynamic potentials. This opens a new pathway for studying black hole phase structures from a purely geometric perspective.
Appendix A: Potential approach
For a spherically symmetric metric
$ds^{2}=-f(r)dt^{2}+\frac{1}{g(r)}dr^{2}+r^{2}d\Omega^{2},$ (47)
the Lagrangian reads
$\mathcal{L}=g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=\sigma.$ (48)
When $\theta=\frac{\pi}{2}$, we obtain the equation of motion
$(\dot{r})^{2}=V_{\text{eff}}(r).$ (49)
The effective potential is given by
$V_{\text{eff}}(r)=g(r)\left[\frac{E^{2}}{f(r)}-\frac{L^{2}}{r^{2}}-\sigma\right],$ (50)
where $E$ and $L$ are defined by Eq.(14). For massless particles in circular orbits, $\sigma=0$, and the unstable circular orbit is denoted as $r_{LR}$. We can derive from
$V_{\text{eff}}(r_{LR})=0,\quad V^{\prime}_{\text{eff}}(r_{LR})=0.$ (51)
This leads to
$f^{\prime}(r_{LR})=\frac{2f(r_{LR})}{r_{LR}}.$ (52)
For massive particles in circular orbits, $\sigma=1$, and the unstable circular orbit is denoted as $r_{0}$. The orbit satisfies
$V_{\text{eff}}(r_{0})=0,\quad V^{\prime}_{\text{eff}}(r_{0})=0.$ (53)
This results in
$L^{2}=\left(\frac{f^{\prime}r^{3}}{2f-rf^{\prime}}\right)_{r=r_{0}},\quad E^{2}=\left(\frac{2f^{2}}{2f-rf^{\prime}}\right)_{r=r_{0}}.$ (54)
Appendix B: Geometric approach
The light ring $r_{LR}$ can also be determined independently using a geometric approach.
Consider the optical metric given in Eq. (7). Its geodesic curvature $\kappa_{g}$ is given by
$\kappa_{g}=\sqrt{\frac{g(r)}{f(r)}}\frac{2f(r)-rf^{\prime}(r)}{2r}.$ (55)
On the light ring $r=r_{LR}$, the geodesic curvature vanishes, yielding
$2f(r_{LR})=r_{LR}f^{\prime}(r_{LR}).$ (56)
This result coincide with Eq. (52).

### v2

@@ IV Critical Exponent from Lyapunov Exponents and Gaussian Curvature
Critical exponents are used to describe the behavior of physical quantities of thermodynamic systems near the critical point and serve as important physical quantities for characterizing phase transitions. At the phase transition point, the small and large black hole branches coexist, leading to discontinuities in both the null Lyapunov exponent and the Gaussian curvature. This section will examine the continuity of the null Lyapunov exponent and Gaussian curvature for the Hayward–Letelier–AdS black hole at the phase transition point $T_{p}$, as well as their discontinuity near $T_{p}$, investigating whether they can serve as order parameters and calculating their critical exponents.
We define the null Lyapunov exponents for the small and large black hole branches within the spinodal region as $\lambda_{S}$ and $\lambda_{L}$, respectively, and the corresponding Gaussian curvatures as $K_{S}$ and $K_{L}$. Their differences are defined as $\Delta\lambda$ and $\Delta K$. When the critical point $T_{p}=T_{c}$, we have $\Delta\lambda=\Delta K=0$. We will fit the numerical coefficients of $\Delta\lambda$ and $\Delta K$ according to the following form Wei and Liu (2018)
$\frac{\Delta\lambda}{\lambda_{c}},\frac{\Delta K}{K_{c}}\sim\alpha(t-1)^{\beta}.$ (30)
Here, $t=T_{p}/T_{c}$, $\alpha$ and $\beta$ are real numbers, the subscript “cc” denotes the value of the Lyapunov exponent and Gaussian curvature at the critical point.
Figure 3: The reduced null Lyapunov exponents and Gaussian curvatures for the coexistence small and large black holes. (a) Null Lyapunov exponents and (b) Gaussian curvature $K_{HL}$ versus reduced temperature $t$ for a Hayward-Letelier-AdS black hole.
As shown in Fig. 3, we plot the normalized differences ${\Delta\lambda}/{\lambda_{c}}$ and ${\Delta K}/{K_{c}}$ as functions of the reduced temperature $t$. When $T_{p}=T_{c}$, ${\Delta\lambda}/{\lambda_{c}}={\Delta K}/{K_{c}}=0$. When $T_{p}$ deviates from $T_{c}$, both ${\Delta\lambda}/{\lambda_{c}}$ and ${\Delta K}/{K_{c}}$ take non-zero values, indicating the emergence of discontinuity. Ultimately, we obtain the following analytical expressions
$\frac{\Delta\lambda}{\lambda_{c}}\simeq 6.9374(t-1)^{0.5761},$ (31)
and
$\frac{\Delta K}{K_{c}}\simeq 49.5398(t-1)^{0.6032}.$ (32)
These results indicate that ${\Delta\lambda}/{\lambda_{c}}$ and ${\Delta K}/{K_{c}}$ can serve as order parameters. However, their critical exponents deviate from the exponent of $1/2$ observed in Reissner-Nordström black holes, suggesting that the geometry and dynamics of regular black holes exhibit richer critical behavior compared to singular black holes.
@@ V Discussion
In this work, we systematically investigate the profound connections between spacetime geometry, thermodynamics, and chaos dynamics during first-order phase transitions of black holes. While the standard thermodynamic methods describe phase transitions through the behavior of thermodynamic potentials, the corresponding evolution of spacetime geometry itself demands a more direct description. Based on this consideration, we examine the possibility of employing the intrinsic geometric quantity Gaussian curvature $K$ as a new probe.
We find that during a first-order phase transition, the Gaussian curvature of unstable null orbits exhibits multivalued behavior inside the spinodal region and coincides precisely with the swallowtail structure of the free energy. This phenomenon can be understood through the known relation $K(r_{LR})=-\lambda^{2}(r_{LR})$, whereby the Gaussian curvature inherits the multivalued nature of the Lyapunov exponent. However, this relation should not be interpreted as a mere translation of the dynamical results into geometric language. The key advance of our work lies in establishing a self-contained geometric diagnostic framework that operates independently of any dynamical calculation. The procedure is purely geometric: first, determining the light ring radius from the geodesic curvature of the optical metric, then computing the corresponding Gaussian curvature via the first fundamental form, and finally, checking for multivalued behavior. This self-consistent procedure yields results in precise agreement with thermodynamic predictions.
@@
Through numerical analysis of the Hayward-Letelier-AdS black hole, we find that when the parameter $\tilde{g}$ enters the spinodal region, the free energy exhibits a swallowtail structure, while both the Gaussian curvature $K$ and the Lyapunov exponent $\lambda$ exhibit multivalued behavior. The temperature intervals over which this occurs align precisely with the thermodynamic spinodal region. Furthermore, we demonstrate that in the absence of a phase transition ($\tilde{g}>\tilde{g}_{c}$), the Gaussian curvature $K$ varies monotonically and the geometric multivaluedness disappears. This confirms that the multivalued behavior of the Gaussian curvature is a geometric signature of the phase transition.
Near the critical point, the scaling laws of the Gaussian curvature and the null Lyapunov exponent for the Hayward-Letelier-AdS black hole deviate from those observed in most black hole solutions, suggesting that regular black holes may exhibit more interesting dynamical and geometric behaviors.
@@
These results advance the paradigm from standard thermodynamic or dynamical frameworks toward a purely geometric perspective. They indicate that black hole first-order phase transitions can be fundamentally understood as a branching behavior in the curvature structure of spacetime. Our work provides the direct evidence of the accompanying degeneracy in spacetime geometry during such phase transitions and demonstrates that Gaussian curvature can serve as a self-consistent diagnostic probe. This opens a new pathway for studying black hole phase structures from a geometric perspective.